OpenStudy (anonymous):

can you help? 4 increased by the product of a number and 3 is less than or equal to -23

OpenStudy (mathstudent55):

Since you need to write an inequality, the first step is to choose a variable to represent the unknown number.

OpenStudy (mathstudent55):

\(\sf 4 ~increased ~by ~the ~product ~of \color{red}{~a ~number} ~and ~3 ~is ~less ~than ~or ~equal ~to ~-23.\) How about we use "x" to represent "a number?"

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

Ok. Now we look at the rest of the sentence, piece by piece, and we build up our inequality.

OpenStudy (mathstudent55):

We have this so far: \(\sf 4 ~increased ~by ~the ~product ~of \color{red}{~a ~number} ~and ~3 ~is ~less \)~than ~or ~equal ~to ~-23. \(~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{red}{x}\)

OpenStudy (mathstudent55):

\(\sf 4 ~increased ~by ~\color{green}{the ~product ~of ~a ~number ~and ~3} ~is ~less \)~than ~or ~equal ~to ~-23. How do we represent the green part now?

OpenStudy (anonymous):

idk what to do with the four i know \[x+3\le23\]

OpenStudy (anonymous):

-23

OpenStudy (mathstudent55):

That is not correct. Stick with me. We will do it one step at a time. Remember that we already picked x for the number. Now we need to express the product of the number and 3. Product is what operation?

OpenStudy (anonymous):

multiply..... i think

OpenStudy (mathstudent55):

Correct. Since product is multiplication, then the product of a number and 3 means 3 times the number. The number is x. What is 3 times x?

OpenStudy (anonymous):

3x

OpenStudy (constanceleak):

4+3n<or = -23?

OpenStudy (mathstudent55):

Good. We have this now: \(\sf 4 ~increased ~by \color{green}{~the ~product ~of ~a ~number} ~and ~3 ~is ~less~than ~or ~equal ~to ~-23.\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{3x}\)

OpenStudy (mathstudent55):

Now we deal with "increased by 4". When you increase a number by another number, you add the number. An increase is an addition.

OpenStudy (anonymous):

thanks you i got it :)

OpenStudy (mathstudent55):

\(\sf \color{purple}{4 ~increased ~by ~the ~product ~of ~a ~number ~and ~3} ~is ~less~than ~or ~equal ~to ~-23.\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{purple}{4 + 3x}\)

OpenStudy (constanceleak):

You welcome : if ur talking to me

OpenStudy (constanceleak):

oh and thx for the medal

OpenStudy (mathstudent55):

Finally we o the end of the sentence. \(\sf \color{purple}{4 ~increased ~by ~the ~product ~of ~a ~number ~and ~3}\color{red}{ ~is ~less~than ~or ~equal ~to ~-23.}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{purple}{4 + 3x} ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{red}{\le -23}\)

OpenStudy (mathstudent55):

The final inequality is: \(3x + 4 \le -23\) If you need to solve it, then you do this: Subtract 4 from both sides. \(3x \le -27\) Divide both sides by 3: \(x \le -9\)

OpenStudy (anonymous):

thanks you are very helpful though i just needed help writing the equation but i fanned you because i appreciate the break down to help me learn

OpenStudy (mathstudent55):

You're welcome. You can use the process I showed for other similar problems. That's why I didn't stop even when you had already been given the answer with no explanation.